OpenStudy (anonymous):

Find the sum of: 46+47+48+49+...+255

OpenStudy (anonymous):

46+ 47+ 48 + 49+... +255+ 255+ 254+ 253+ 252+ .. +46= 301+301+301 + 301 + ... +301=2 S (255-46+1)301= 2 S\\ How much is S?

OpenStudy (anonymous):

\[ S= \frac {(255-46+1)301}{2}=\frac{63210}{2}=31605 \]